OpenStudy (anonymous):

Which of the following points lies on the graph of y = x^2 - 2x + 6? A.) (-3, 21) B.) (-2, 0) C.) (-1, 10)

OpenStudy (anonymous):

the answer is A

mathslover (mathslover):

You can use hit and trial method. Plug-in the values of x and y from the options. I will do one for you as an example : In option B) We have (-2,0) So, x = -2 and y = 0 Pluggin in these values in the given equation : \(\sf{y^2 = x^2 - 2x + 6}\\ \textbf{Since, y = 0 and x= -2 } \\ \sf{(0)^2 = (-2)^2 - 2(0) + 6\\ 0 = 4 - 0 + 6 \\ 0 = 10}\\ \textbf{This is not possible as 10 is not equal to 0}\\ \) Therefore option B is incorrect

OpenStudy (anonymous):

Will you help me with a few more questions, please? If you don't mind.

mathslover (mathslover):

@mimipink16 : Yes it is A. But instead of giving direct answers you must direct the user (asker) to the correct solution. If you will just provide an answer then how will the user know about the solution or the way to get to the answer. So, please in future don't repeat this mistake. Also, OpenStudy 's aim is to help users not to cheat them. Refer OpenStudy Code-of-Conduct for future referrence : http://openstudy.com/code-of-conduct

mathslover (mathslover):

I will @xo_kansasprincess_xo , I am always here to help those users who help themselves first :) Well first let me know , did you get this one?